# RetinexDualV2: Physically-Grounded Dual Retinex for Generalized UHD Image Restoration

### 30 Mar 2026 v1

@@ 5 Conclusion
We presented RetinexDualV2, a unified and parameter-efficient framework for diverse UHD image restoration tasks. By integrating a Task-Specific Physical Grounding Module (TS-PGM) and Physical-conditioned Multi-head Self attention (PC-MSA), our model explicitly leverages degradation-aware physical priors to guide a dual-branch Retinex decomposition network. This design gracefully overcomes the generalizability limits of purely data-driven black-box architectures, adapting seamlessly to various degradations without task-specific modifications. RetinexDualV2’s robustness is evidenced by its competitive performance, notably securing 4th place in the NTIRE 2026 Day and Night Raindrop Removal Challenge and 6th place in the Joint Noise Low-light Enhancement Challenge. Our extensive evaluations underscore the value of integrating physical priors, highlighting meaningful progress toward interpretable foundation models for adverse weather image restoration.

### 09 Apr 2026 v2

@@ 5 Conclusion
We presented RetinexDualV2, a unified and parameter-efficient framework for diverse UHD image restoration tasks. By integrating a Task-Specific Physical Grounding Module (TS-PGM) and Physical-conditioned Multi-head Self-Attention (PC-MSA), our model explicitly leverages degradation-aware physical priors to guide a dual-branch Retinex decomposition network. This design gracefully overcomes the generalizability limits of purely data-driven black-box architectures, adapting seamlessly to various degradations without task-specific modifications. RetinexDualV2’s robustness is evidenced by its competitive performance, notably securing 4th place in the NTIRE 2026 Day and Night Raindrop Removal Challenge and 5th place in the Joint Noise Low-light Enhancement Challenge. Our extensive evaluations underscore the value of integrating physical priors, highlighting meaningful progress toward interpretable foundation models for adverse weather image restoration.